# HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models

### v1

@@ Abstract
The ubiquitousness of social media has led to the need for reliable and efficient detection of offensive content to limit harmful effects. This has led to a proliferation of datasets and models related to detecting offensive content. While sophisticated models have attained strong performance on individual datasets, these models often do not generalize due to differences between how “offensive content” is conceptualized, and the resulting differences in how these datasets are labeled. In this paper, we introduce HateCOT, a dataset of 52,000 samples drawn from diverse existing sources with explanations generated by GPT-3.5-Turbo and human-curated. We show that pre-training models for the detection of offensive content on HateCOT significantly boots open-sourced Language Models on three benchmark datasets in both zero and few-shot settings, despite differences in domain and task. We further find that HateCOT enables effective K-shot fine-tuning in the low-resource settings. Our repository is available at https://github.com/hnghiem-usc/hatecot.
@@ 1 Introduction
Compounding the challenge, reliable detection of offensive content requires significant data resources. Sophisticated models tend to be data-hungry, and the process of curating a dataset tailored to specific use cases can be costly, time-consuming, and emotionally challenging for annotators Founta et al. (2018); Toraman et al. (2022). The typical pipeline consists of collecting samples based on topic-relevant key words, then recruiting either crowdworkers or experts to annotate data before developing classification models Paullada et al. (2021). Each step incurs investment and may inject subtle idiosyncrasies to the downstream dataset, further limiting potential transferable usefulness to related tasks Fortuna et al. (2021).
@@ 2.2 LLMs in Offensive Speech Classification
Pre-trained language models have exhibited remarkable ability in text generation Celikyilmaz et al. (2020). Recent scaled-up LLMs such as the GPT-3 and later models are capable of astoundingly fluent, convincing and knowledge-infused outputs Zhang et al. (2023). LLMs with hundreds or more parameters even exhibit reasoning capabilities Wei et al. (2022, 2021), leading to a flurry of research on prompting techniques to harness their prowess, such as Chain-of-thought (COT), Tree-of-thought etc. (Yao et al., 2023; Diao et al., 2023). An interesting line of research leverages LLMs to efficiently generate high volumes of synthetic data for tasks with training resource is scarce Puri et al. (2020); Bao et al. (2023); Whitehouse et al. (2023). We build upon these works to construct a dataset that can induce smaller LLMs to efficiently adapt to new categories of offensive content by leveraging their provided definitions.
@@ 3.1 Data Selection
    Definition: each dataset should have the associated definitions with each label available. This criteria is important to generate informative explanations.
@@
    Target / Rationale: the dataset should elicit the targets and/or reasoning to justify their choice of label (rationale) of offensive sentiments, whether through discrete labels or free text.
@@
Salminen et al. (2018)
137,098
5, 418
Y, F
Hateful, Neutral
Qian et al. (2019)
34,000
5, 034
G
Not Hate, Hate
Sap et al. (2019)
44,671
6, 033
G, R, T
Not Offensive, Offensive
Vidgen et al. (2021a)
27,494
6, 717
R
Neutral, Person Directed Abuse, Affiliation Directed Abuse, Identity Directed Abuse
Vidgen et al. (2021b)
10,152
7, 209
S
None, Derogation, Dehumanization, Animosity, Support, Threatening
Sachdeva et al. (2022)
135,556
7, 272
Y, T, R
Not Hate speech, Hate Speech
Hartvigsen et al. (2022)
274,186
7, 239
S
Benign, Toxic
Toraman et al. (2022)
100,000
7, 215
T
Normal, Offensive, Hate
Total
52, 137
(a) Datases used to create training corpus. Sample Size denotes the number of chosen samples from corresponding dataset included in the training corpus.
@@
HateCheck
3,728
50:50
300
500
S
Non-hateful, Hateful
HateXplain
20,148
60:40
200
400
G, T
Normal, Offensive, Hate
Latent_Hate
19,112
60:40
200
400
T
Not Hate, Explicit Hate, Implicit Hate
Implicit_Hate
4,153
60:40
150
T
White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation
@@
Yang et al. (2023) showed that using GPT-3 to augment human-written rationales creates linguistically coherent while maintains faithfulness to the original content. We use the template shown in Figure 1 to generate the explanation for each input guided by the available annotations on label, target, and rationale from the chosen corpora. For datasets that contains multiple annotations for each sample, we select the ultimate label via majority voting and concatenate individual annotations on the targets and/or rationales into a single string separated by the delimiter "|".
@@ 3.2.1 Setup
Figure 4 illustrates the Alpaca-styled template to format each post with its corresponding label, generated explanation, definitions along with the instruction into blocks of an input prompt (Taori et al., 2023). Using the described corpus, we supervised finetune the open-sourced model Llama 2 Chat-HF of 7 billion parameters (hereby referred to as Llama 7B) (Touvron et al., 2023) via LoRA techniques (Technical details specified in Appendix A.2) (Hu et al., 2021). Then, we perform zero-shot classification using the same template to prompt the fine-tuned models to generate the explanation and label for posts drawn from the 3 test datasets. For this part, posts are drawn using K-shot sampling (equal number of samples for each class) on the Validation set of the test data (K val in Table 1(b)). We omit Implicit_Hate due to this set’s markedly different 6-label space.
@@
For the first experiment on the ratio of neutral vs. non-neutral categories, the training data is split into subsets with their respective distribution between the neutral (NE) class and non-neutral class(es) (NN) described by the following formula: $NN=R*NE$, where $R\in\{1,2,3,4\}$ is the ratio coefficient. We set the number of explanations per sample to 2 for this part. For the second experiment, we construct the subsets by varying the degree of reasoning diversity $D\in\{1,2,3,4\}$, of each post.
@@
We extract the explanation and predicted labels after their respective tag in the generated output. If the models generate multiple items from the dataset’s label space, we select the first legal label. If no acceptable output is obtained, we randomly select an item in the label space.
@@ 3.2.2 Insights and Augmentation
We report the model’s macro F1-scores on the validation set of each configuration in Table 2. A balanced distribution between neutral and non-neutral classes in the training corpus is beneficial, as reflected by the substantially high mean F1-score of 0.643 when R=1. On the other hand, having 3 explanations per sample (D=3) achieves the best overall performance across 3 test sets, consistent with Ho et al. (2022)’s findings on the benefit of multiple reasoning paths. Nevertheless, performance markedly degrades when D=4. Based on manual analysis, we hypothesize that the quality of generated outputs degrades as the degree of diversity increases, consequently affecting the performance of models trained on this data. These insights are valuable to practitioners when curating synthetic data from LLMs.
From these results, we augment the training corpus by selecting an additional approximately 1,800 samples from each of the 8 datasets while preserving the 1:1 balanced ratio of neutral to non-neutral classes. Then we collect 3 explanations per sample using the described mechanism, resulting in a final corpus of 52,137 samples (Table 0(a)), hereby referred to as the HateCOT corpus.
@@ 4.1 Zero-shot Classification
This setting involves prompting the models to perform the classification task with no in-context examples via 2 modes: No Explanation, which prompts the model to directly predict the label for the given input, and With Explanation, which requires a justification before the predicted label. We finetune the base models using only HateCOT and evaluate their performance on the 3 test sets as in Section 3.2.2. Relevant technical details on pre-training are presented in Appendix A.2.
With results presented in Figure 1(a), we observe that the smaller Flan-T_L and OPT-IML are not able to generate explanations even when prompted to. In contrast, their scaled-up counterparts could follow instructions at all settings, indicating that size matters to effectively exploit instruction-tuning Wei et al. (2021)). Asking Base (off-the-shelve) models to provide explanations before label provides observable boost to Llama models in HateCheck and HateXPlain datasets, but actually hampers performance in Latent_Hate (and its derivative Implicit_Hate), a dataset noted for to be challenging for classifiers due to its covert nature ElSherief et al. (2021b).
Finetuning on HateCOT unanimously enables all models to generate explanations. While smaller models receive no observable boost in their performance, larger models (COT-T5-XL, Llama 7B, Llama 13B) are considerably enhanced compared to their base counterparts. For COT-T5-XL, with the exception of HateXplain, HateCOT-trained version attain F1 scores increment of 7.6% (HateXplain and 9.5% (Latent_Hate) over the base counterpart without explanations. Llama 7B observes 23.9%, 25.6%, 10% increment for HateCheck, HateXplain, Latent_Hate, respectively. These statistics are 27.9%, 118.5%, and 10.2% for Llama 13B. For Implicit_Hate, all models struggle to attain satisfactory performance, with COT-T5-XL leading the pack. Llama models’ performance improve with HateCOT, but still not much better than chance.
@@
(b) Macro F-1 scores for models in zero-shot setting with explanation after K-shot in-domain finetuning at various values of K. Dashed line represents finetuned base models, solid line represents models pre-trained on HateCOT, then in-domain finetuned.
@@ 4.2 In-Domain Finetuning
We further finetune COT-T5-XL, Llama 7B, Llama 13B using data from the training set of the test datasets, including the Implicit_Hate set with 6-way classification. To simulate low-resource settings, we select 256 samples uniformly at random from each class of labels from the datasets, then augment them with explanations in the same fashion as described in Section 3.1. We perform K-shot sample (K $\in\{32,64,128,256\}$) to finetune and evaluate the models on the test sets. This process applies to both Base and Pretrained versions.
@@ 4.3 In-context Learning
As an alternative to in-domain fine-tuning, we investigate the models’ performance using in-context learning (ICL), when a number of complete examples are provided as part of the input prompt. We select 1 sample from each class in the training data of each dataset, then obtain its the associated explanation from GPT-3.5-Turbo. The sets of post, explanation and label are arranged in the same format in the same template as in Figure 4. We run inference for classification results over 5 random seeds, which also randomly permutes the samples’ order.
@@
4.4 Overall Assessment
Our experiments demonstrate that HateCOT confers various benefits relating to generalizing offensive speech detection. Priming LLMs using HateCOT via supervised training significantly enhances zero-shot predictive power. Competitive performance is possible to obtained using significantly less in-domain training data, a bottleneck factor on resources in the traditional pipeline. Our findings suggest suggest a cost-efficient alternative approach: practitioners could obtain human-derived annotation and rationales for K-shot samples, then augment them to acquire explanations using LLMs as outlined. This process may be iterated until desired performance is reached according to some guiding criteria, minimizing the waste of resources.
@@ 5 Quality of Explanations
We analyze the quality of the explanations generated by various versions of 3 models COT-T5-XL, Llama 7B, Llama 13B. Both recent literature Liu et al. (2023); Li et al. (2023) and our validation experiment (reported in QA 5.1 below) corroborate OpenAI’s GPT-4’s high agreement with human annotators. Thus, we leverage GPT-4’s judgement to compare our models’ outputs.
@@
We randomly select 500 samples per model whose explanations from both the Base and HateCOT-pretrained versions agree on the predicted label from all combined test sets. Using the template in Figure 5, we ask GPT-4 to select its preferred explanation for the purpose of content moderation based on fluency, soundness and the alignment with the definition of the chosen label. We present the explanations from these 2 versions in a randomized order for the same post. From this pool, we recruit a human annotator to perform the same labelling task on 150 random samples. Human decisions match GPT-4’s 68% of the times, indicating considerable agreement (Appendix A.4) .
Converting GPT-4’s ratings to an ordinal scale (0 if Base model is preferred, 1 otherwise), we perform the Wilcoxon rank sum test with the null hypothesis $H_{0}$: there is no significant difference between the number of times Pretrained or Base model is preferred over the other. Table 3 shows the overwhelming preference for explanations of the Pretrained models. That $p$-value $\ll$ $\alpha=0.05$ rejects the null hypothesis and provides evidence of HateCOT-induced improvement on the quality of explanation, particularly for COT-T5-XL.
Model Base Pretrained p-value
COT-T5-XL
27   (5.4%)
473 (94.6%)
Llama 7B
109 (21.8%)
391 (78.2%)
Llama 13B
122 (24.4%)
378 (75.6%)
Table 3: Number of times outputs of Base vs Pretrained models are preferred by GPT-4 across 500 samples. *** indicates $p$-value $\ll$ 0.05 by the Wilcoxon rank test.

### v2

@@ Abstract
Warning: This paper contains examples of very offensive content. The ubiquitousness of social media has led to the need for reliable and efficient detection of offensive content to limit harmful effects, leading to the proliferation of datasets and models related to detecting offensive content. While sophisticated models have attained strong performance on individual datasets, they often do not generalize due to differences between how “offensive content” is conceptualized, and how these datasets are labeled. In this paper, we introduce HateCOT, a dataset of 52,000 samples drawn from diverse existing sources with explanations generated by GPT-3.5-Turbo and human-curated. We show that pre-training models for the detection of offensive content on HateCOT significantly boots open-sourced Language Models on three benchmark datasets in both zero and few-shot settings, despite differences in domain and task. We further find that HateCOT enables effective K-shot fine-tuning in low-resource settings.111 Our repository is available at [REDACTED]
@@ 1 Introduction
Compounding the challenge, reliable detection of offensive content requires significant data resources. Sophisticated models tend to be data-hungry, and the process of curating a dataset tailored to specific use cases can be costly, time-consuming, and emotionally challenging for annotators Founta et al. (2018); Toraman et al. (2022). The typical pipeline consists of collecting samples based on topic-relevant key words, then recruiting either crowdworkers or experts to annotate data before developing classification models Paullada et al. (2021). Each step incurs investment and may inject subtle idiosyncrasies proportionate to the size of the downstream dataset, further limiting potential transferable usefulness to related tasks Fortuna et al. (2021). Furthermore, the size of a dataset does not necessarily guarantee cross-main transferrability Poletto et al. (2021); Fortuna et al. (2021).
@@ 2.2 LLMs in Offensive Speech Classification
Pre-trained language models have exhibited remarkable ability in text generation Celikyilmaz et al. (2020). Recent large-size LLMs such as GPT-3 and later models are capable of astoundingly fluent, convincing and knowledge-infused outputs Zhang et al. (2023). LLMs with hundreds or more parameters even exhibit reasoning capabilities Wei et al. (2022, 2021), leading to a flurry of research on prompting techniques to harness their prowess, such as Chain-of-thought (COT), Tree-of-thought etc. (Yao et al., 2023; Diao et al., 2023). An interesting line of research leverages LLMs to efficiently generate high volumes of synthetic data for tasks with training resource is scarce Puri et al. (2020); Bao et al. (2023); Whitehouse et al. (2023). We build upon these works to construct a dataset that can induce smaller LLMs to efficiently adapt to new categories of offensive content by leveraging their provided definitions.
@@ 3.1 Data Selection
    Definition: each dataset should have the associated definitions with each label available (Figure 6). This criteria is important to generate informative explanations.
@@
    Target / Rationale: the dataset should elicit the a) targets and/or b) rationales, which are fragments of free texts by human annotators to convey some understanding of the corresponding post. Figure 4 shows several examples of rationales, demonstrating their unsuitability to be used as explanations in their native format.
@@
Salminen et al. (2018) 137,098 5, 418 Y, F Hateful, Neutral
Qian et al. (2019) 34,000 5, 034 G Not Hate, Hate
Sap et al. (2019) 44,671 6, 033 G, R, T Not Offensive, Offensive
Vidgen et al. (2021a) 27,494 6, 717 R Neutral, Person Directed Abuse, Affiliation Directed Abuse, Identity Directed Abuse
Vidgen et al. (2021b) 10,152 7, 209 S None, Derogation, Dehumanization, Animosity, Support, Threatening
Sachdeva et al. (2022) 135,556 7, 272 Y, T, R Not Hate speech, Hate Speech
Hartvigsen et al. (2022) 274,186 7, 239 S Benign, Toxic
Toraman et al. (2022) 100,000 7, 215 T Normal, Offensive, Hate
Total - 52, 137
(a) Datasets used to create training corpus. Sample Size denotes the number of chosen samples from corresponding dataset included in the training corpus.
@@
HateCheck 3,728 50:50 300 500 S Non-hateful, Hateful
HateXplain 20,148 60:40 200 400 G, T Normal, Offensive, Hate
Latent_Hate 19,112 60:40 200 400 T Not Hate, Explicit Hate, Implicit Hate
Implicit_Hate 4,153 60:40 - 150 T White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation
@@
Yang et al. (2023) showed that using GPT-3 to augment human-written rationales creates linguistically coherent while maintains faithfulness to the original content. We use the template shown in Figure 1 to generate the explanation for each input guided by the available annotations on label, target, and rationale from the chosen corpora. For datasets that contain multiple annotations per sample, we select the ultimate label via majority voting and concatenate annotations on the targets and/or rationales into a single string delimited by "|".
@@ 3.2.1 Setup
Figure 5 illustrates the Alpaca-styled template to format each post with its corresponding label, generated explanation, definitions along with the instruction into blocks of an input prompt (Taori et al., 2023). Using the described corpus, we supervised finetune the open-sourced model Llama 2 Chat-HF of 7 billion parameters (hereby referred to as Llama 7B) (Touvron et al., 2023) via LoRA techniques (Technical details specified in Appendix A.2) (Hu et al., 2021). This model is selected as the primary representative for this stage due to its manageable size and strong classification performance Touvron et al. (2023). Then, we perform zero-shot classification using the same template to prompt the fine-tuned models to generate the explanation and label for posts drawn from the 3 test datasets. For this part, posts are drawn using K-shot sampling (equal number of samples for each class) on the Validation set of the test data (K val in Table 2(b)). We omit Implicit_Hate at this stage due to this set’s markedly different 6-label space.
@@
For the first experiment on the ratio of neutral vs. non-neutral categories, the training data is split into subsets with their respective distribution between the neutral (NE) class and non-neutral class(es) (NN) described by the following formula: $NN=R*NE$, where $R\in\{1,2,3,4\}$ is the ratio coefficient. We set the number of explanations per sample to 2 , the smallest value that still enables the benefit of reasoning diversity. For the second experiment, we construct the subsets by varying the degree of reasoning diversity $D\in\{1,2,3,4\}$, of each post.
@@
We extract the explanation and predicted labels after their respective tag in the generated output. If the models generate multiple items from the dataset’s label space, we select the first admissible label. If no acceptable output is obtained, we randomly select an item in the label space.
@@ 3.2.2 Insights and Augmentation
We report the model’s macro F1-scores on the validation set of each configuration in Table 2. A balanced distribution between neutral and non-neutral classes in the training corpus is beneficial, as reflected by the substantially high mean F1-score of 0.643 when R=1. On the other hand, having 3 explanations per sample (D=3) achieves the best overall performance across 3 test sets, consistent with Ho et al. (2022)’s findings on the benefit of multiple reasoning paths. Nevertheless, performance markedly degrades when D=4. Based on manual analysis, we hypothesize that the quality of generated outputs degrades as the degree of diversity increases, consequently affecting the performance of models trained on this data. From these results, we augment the training corpus by selecting an additional approximately 1,800 samples from each of the 8 datasets while preserving the 1:1 balanced ratio of neutral to non-neutral classes. Then we collect 3 explanations per sample using the described mechanism, resulting in a final corpus of 52,137 samples (Table 1(a)), hereby referred to as HateCOT.
@@ 4.1 Zero-shot Classification
This setting prompts models to perform the classification task with no in-context examples via 2 modes: No Explanation, which prompts the model to directly predict the label for the given input, and With Explanation, which requires a justification before the predicted label. We finetune the base models using only HateCOT and evaluate their performance on the 4 test sets as in Section 3.2.2 (technical details in Appendix A.2).
With results presented in Figure 2(a), we observe that the smaller Flan-T5-L and OPT-IML are unable to generate explanations when prompted. In contrast, their scaled-up counterparts could follow instructions at all settings. Asking Base (off-the-shelve) models to provide explanations before label provides observable boost to Llama models in HateCheck and HateXPlain datasets, but actually hampers performance in Latent_Hate and its derivative Implicit_Hate, datasets notably challenging due to its covert nature ElSherief et al. (2021b).
Model Choice Matters
Finetuning on HateCOT unanimously enables all models to generate explanations. While smaller models receive no observable boost in their performance, larger models (COT-T5-XL, Llama 7B, Llama 13B) are considerably enhanced compared to their base counterparts. For COT-T5-XL, with the exception of HateXplain, HateCOT-trained version attain F1 scores increment of 7.6% (HateXplain and 9.5% (Latent_Hate) over the base counterpart without explanations. Llama 7B observes 23.9%, 25.6%, 10% increment for HateCheck, HateXplain, Latent_Hate, respectively. These statistics are 27.9%, 118.5%, and 10.2% for Llama 13B. For Implicit_Hate, all models struggle to attain satisfactory performance. Llama models’ performance improve with HateCOT, but still not much better than chance.
The reduction in performance of HateCOT-pretrained models compared to their Base counterparts without explanations (Flan-T5-L and OPT-IML on almost all test sets, COT-T5-XL on HateXPlain and Implicit_Hate) is in line with literature as COT-style prompting tends to favor larger models Wei et al. (2022); Wang et al. (2024); Suzgun et al. (2023). These models’ impacted performance when asked for explanation is unsurprising considering their significantly smaller size and original pre-training data, especially in zero-shot settings. Iyer et al. (2022); Chung et al. (2022).
@@
(b) Macro F-1 scores for models in zero-shot setting with explanation after K-shot in-domain finetuning at various values of K. Dashed line represents finetuned base models, solid line represents models pre-trained on HateCOT, then in-domain finetuned. For each dataset, the horizontal dashed line represents the base version of the best performing model at K=256 which is finetuned using the entire training data without any rationale for comparison, denoted as Best model (Full data, No. exp).
@@ 4.2 In-Domain Finetuning
We further finetune COT-T5-XL, Llama 7B, Llama 13B using data from the training portion of the test datasets, including the Implicit_Hate set. To simulate low-resource settings, we choose 256 samples uniformly at random from each class, then augment them with explanations as described in Section 3.1. We perform K-shot sample (K $\in\{32,64,128,256\}$) to finetune both Base and Pretrained versions of the models.
@@
We further select the best performing model at K=256 for each dataset and finetune their Base versions with the entire training data, without explanations and definitions. As shown in Figure 2(b) and Table 4, in-domain finetuning with HateCOT enables models to attain competitive results even with only a fraction of the training data. Furthermore, prompting for explanations enables Llama 13B and COT-T5-XL to attain performance surpassing using the full training data for Latent_Hate and Implicit_Hate.
@@ 4.3 In-context Learning
As an alternative to in-domain fine-tuning, we investigate the models’ performance using in-context learning (ICL), when a number of complete examples are provided as part of the input prompt. We select 1 sample from each class in the training data of each dataset, then obtain its the associated explanation from GPT-3.5-Turbo. The sets of post, explanation and label are arranged in the same format in the same template as in Figure 5. We run inference for classification results over 5 random seeds, which also randomly permutes their order.
@@
4.4 Assessment and Recommendations
From empirical observations, we make the following recommendations to construct a cost-efficient pipeline for classifier on novel domains:
  • $\triangleright$
    Practitioners should choose model of sufficient number of parameters for the task. Larger instruction-tuned LLMs appear to more effectively capitalize on HateCOT pretraining regimen before in-domain finetuning.
  • $\triangleright$
    Instead of devoting resources to curate substantial training data, practitioners could focus on obtaining high quality annotations for representative rationales, and augment them into explanations using their LLM of choice. Alternatively, practitioners may choose to curate the explanations organically to achieve certain desired thematic qualities. This process may be iterated until targeted performance is reached according to some guiding metrics with acceptable quality of explanation.
GPT-4 Human (Frequency Count) Human (Majority Vote)
Model Base Pretrained $p$-val. Base Pretrained Base Pretrained
COT-T5-XL 27 (5.4%) 473 (94.6%) *** 62 (24.8%) 188 (75.2%) 6 (12%) 44 (88%)
Llama 7B 109 (21.8%) 391 (78.2%) *** 109 (43.6%) 141 (56%) 19 (38%) 31 (62%)
Llama 13B 122 (24.4%) 378 (75.6%) *** 114 (45.6%) 136 (54.4%) 22 (44%) 28 (56%)
Table 3: Comparison of Base and Pretrained models in GPT-4 and Human Evaluation. *** indicates $p$-value $\ll$ 0.05 by the Wilcoxon rank test. Output preferred by at least 3 out of 5 annotator wins the Majority Vote for that sample.
@@ 5 Quality of Explanations
We analyze the quality of the explanations generated by various versions of 3 models COT-T5-XL, Llama 7B, Llama 13B through evaluation by both human annotators and OpenAI’s GPT-4 following similar literature Liu et al. (2023); Li et al. (2023). An additional assessment of Target Identification is presented in Appendix A.5.
@@
We randomly select 500 samples per model whose explanations from both the Base and HateCOT-pretrained versions agree on the predicted label from all combined test sets. Using the template in Figure 6, we ask GPT-4 to select its preferred explanation for the purpose of content moderation based on fluency, soundness and the alignment with the definition of the chosen label. We present the explanations from these 2 versions in a randomized order for the same post. Using a similar setup, we recruit 13 annotators from the crowdsource platform Amazon Mechanical Turk and solicit their rating on 50 samples per model (Appendix A.4). Each post is annotated by 5 humans, resulting in 750 annotations across 3 models.
Converting GPT-4’s ratings to an ordinal scale (0 if Base model is preferred, 1 otherwise), we perform the Wilcoxon rank sum test with the null hypothesis $H_{0}$ (no significant difference between the number of times Pretrained or Base model is preferred over the other). In Table 3, we observe the preference for explanations of the Pretrained models across all settings, even when considering the human tally by using majority vote ($\geq$ 3 out of 5 annotations). The human preference margin is smaller for Llama 13B Pretrained, likely due to this model’s already strong generative capabilities.